Today, interest rates on 1-year T-bonds yield 1.3%, interest rates on 2-year T-bonds yield 2.4%, and interest rates on 3-year T-bonds yield 3.4%.

a. If the pure expectations theory is correct, what is the yield on 1-year T-bonds one year from now? Be sure to use a geometric average in your calculations. Do not round intermediate calculations. Round your answer to four decimal places.

   %

b. If the pure expectations theory is correct, what is the yield on 2-year T-bonds one year from now? Be sure to use a geometric average in your calculations. Do not round intermediate calculations. Round your answer to four decimal places.

   %

c. If the pure expectations theory is correct, what is the yield on 1-year T-bonds two years from now? Be sure to use a geometric average in your calculations. Do not round intermediate calculations. Round your answer to four decimal places.

   %
Answer #1

One year yield is 1.3%.

Two year yield is 2.4%.

Three year yield is 3.4%.

a.

(Ef + 1) = (h+1) (Th+I)

(1 + 0.013) x (1 + 91.1) = (1 +0.024)

(1+ 41,1) = 1.048576 1.013

(1+1.11 = 1.035119

Y1,1 = 0.035119

Yield on 1 year bond , one year from now is 3.5119%.

b.

(1+yı) x (1 +41,2) = (1 +43)

(1 + 0.013) x (1 + 91.2)2 = (1 + 0.034)

(1 + 1.1.1.105507304 1.013

(1 + 1.2)= 1.1.0913201422

Y1.2 = 1.044663 - 1

Y1.2 = 0.044663

Yield on 2 year bond , one year from now is 4.4663%.

c.

(1 + y2)2 x (1 + 42,1) = (1 +43)

(1 +0.024)2 (1 + 421) = (1 + 0.034)

(1+21) = 1.10507304 1.048576

(1+21) = 1.054294

41,1 = 0.054294

Yield on 1 year bond , two year from now is 5.4294%.